# Feedback Delay Network Optimization

### v1

@@ 7 Conclusion
Favorable results were obtained in this study with a DiffFDN employing as few as four delay lines. In objective evaluations, we demonstrated that the proposed method reduces the width of the modal excitation distribution, thereby decreasing the prominence of the loudest modes. Listening test results further confirmed the success of this approach in attenuating artifacts commonly associated with undesired coloration, making the DiffFDN close to an ideally colorless, smooth reverberator.
@@
The PyTorch implementation of the proposed method can be found on the online repository  [^1] https://github.com/gdalsanto/diff-fdn-colorless/. Listening test stimuli examples can be found on the dedicated website  [^2] http://research.spa.aalto.fi/publications/papers/eurasip-colorless-fdn/.

### v2

@@ 7 Conclusion
Favorable results were obtained in this study with a DiffFDN employing as few as four delay lines. In objective evaluations, we demonstrated that the proposed method reduces the width of the modal excitation distribution, thereby decreasing the prominence of the loudest modes. Listening test results further confirmed the success of this approach in attenuating artifacts commonly associated with undesired coloration, making the DiffFDN close to an ideally colorless, smooth late reverberation.
@@
The PyTorch implementation of the proposed method, together with the code used to generate the figures, can be found on the online repository  [^1] https://github.com/gdalsanto/diff-fdn-colorless/. Listening test stimuli examples can be found on the dedicated website  [^2] http://research.spa.aalto.fi/publications/papers/eurasip-colorless-fdn/.